Mrs. Brook lacks 7 years from being five times as old as her son. Six years from now she will lack 3 years from being three times as old as her son is then. Find each of their ages.
Answer by lwsshak3(11628)

Mrs. Brook lacks 7 years from being five times as old as her son. Six years from now she will lack 3 years from being three times as old as her son is then. Find each of their ages.
**
At present:
let x=son's age
(5x-7)=mother's age (mother is 7 yrs younger than 5 times son's age)
..
6 years hence:
(x+6)=son's age
(5x-7)+6=mothers age
..
(5x-7)+6=3(x+6)-3 (mother is 3 yrs younger than 3 times son's age)
5x-1=3x+18-3=3x+15
2x=16
x=8
5x-7=33
ans:
son's age=8
mother's age=33